OpenStudy (anonymous):

Integrate (x-3/2x^-1/2 / -1/2

OpenStudy (raden):

is this your question ? 

OpenStudy (anonymous):

No Wait a sec I will try to draw it

OpenStudy (anonymous):

I'm seeing this \[-\frac{x-\frac{ 3 }{ 2 }x^\frac{ -1 }{ 2 }}{\frac{ 1 }{ 2 }}\] ???? yes please draw it :D

OpenStudy (anonymous):

The image you have above is correct Phaen

OpenStudy (anonymous):